MILLIMETERWAVE AND FOURIER-TRANSFORM EMISSION SPECTRA OF THE BiS RADICAL

K. IZUMI, K. KAWAGUCHI, Department of Chemistry, Okayama University, Okayama, Japan, 700-8530; \underline{E. A. COHEN}, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, California 91109-8099; K. D. SETZER, and E. H. FINK, Physikalische Chemie-Fachbereich C, Bergische Universität Wuppertal, D-42097 Wuppertal, Germany.

The millimeterwave rotational spectrum of BiS in its X_{1} ^{2}Π_{1/2} state was observed in the frequency range of 150-300 GHz. BiS was produced in a high-temperature oven by a discharge in a mixture of Bi vapor and CS2. A White-type multipath cell was used to enhance the absorption pathlength. Near infrared bands of the transition were measured by Fouriertransform emission spectroscopy in the 6400-7400 cm^{−1} region, where BiS was produced by reaction of bismuth and sulfur vapor and excited by collisional energy transfer from the metastable a^{1}Δ_{g} electronic state of O_{2}. A simultaneous analysis of millimeterwave and FT data was carried out to give rotational, fine and hyperfine constants for the X_{1} ^{2}Π_{1/2} and X_{2} ^{2}Π_{3/2} states. Ninety seven rotational ΔJ = 1 features from J′ = 23.5 to 41.5 and 545 NIR features representing assignments of a wide range of J’s were included in the fit. The hyperfine parameters are consistent with those of BiO^{a} with slightly less unpaired electron density in the antibonding π orbital on the Bi atom. Examples of the spectra will be shown. The fitting procedure with spcat^{b} and the resulting parameters will be discussed.

aE. A. Cohen, D. M. Goodridge, K. Kawaguchi, E. H. Fink, and K. D. Setzer, J. Mol. Spectrosc. 239 (2006) 16-23.

bH. M. Pickett, J. Mol. Spectrosc. 148 (1991) 271-377.